OpenStudy (hesan):

1 nm = ? mm As 1m =10^-6 nm 10^-3 mm = 10^-6 nm 10^-3 /10^-6 mm = 1 nm i.e . 1nm = 1000mm but it is not so.. why

ganeshie8 (ganeshie8):

nano... you must remember it cuz nano technology is all around

ganeshie8 (ganeshie8):

\(nano = 10^{-9}\)

OpenStudy (hesan):

1 um = ? mm As 1m =10^-6 um 10^-3 mm = 10^-6 um 10^-3 /10^-6 mm = 1 um i.e . 1um = 1000mm but it is not so.. why

OpenStudy (hesan):

sorry that was micro meter

OpenStudy (unklerhaukus):

[um] = 10^-6 [m]

ganeshie8 (ganeshie8):

\(\large 1um = 1um \times \frac{10^{-3} mm}{1 um}\)

OpenStudy (hesan):

if we do it so with time then 1 hr = 60 min 3600 sec = 60 min 1sec = 60/3600 min 1 sec = 1/60 min that gives the correct result..

OpenStudy (unklerhaukus):

[m] = 10^6 [um]

OpenStudy (unklerhaukus):

[mm] = 10^-3 [m]

OpenStudy (hesan):

@UnkleRhaukus 1m = 10^-6 not 10^6 m 10^ 6 is for mega

OpenStudy (unklerhaukus):

micro meters are tiny , there are i million (ie 10^6) micrometers in one meter

ganeshie8 (ganeshie8):

1 meter = 10^6 micro meter

OpenStudy (hesan):

@UnkleRhaukus yes u are right thats the mistake.. thanks

OpenStudy (unklerhaukus):

1 [um] = ? [mm] 10^-6 [m] = ? 10^-3 [m] 10^-6 [m]/10^-3 [m] = ?